OpenStudy (anonymous):

Mrs. Grijalva hired a landscaping company to landscape her back yard. The company has a standing policy to plant 3 bushes for every tree planted. The company charges $75 for each tree planted and $20 for each bush. Mrs. Grijalva was billed for $810 after the planting was complete. How many bushes did the company plant in Mrs. Grijalva’s back yard? A. 2 B. 6 C. 18 D. 54

OpenStudy (jdoe0001):

so she got billed 810 they planted "t amount" of trees they also planted "b amount" of bushes so her invoice was => (t * $75) + (b * $20) = $810

OpenStudy (jdoe0001):

what equation would that give us?

OpenStudy (anonymous):

75+20=810?

OpenStudy (jdoe0001):

well, you're missing the variables but close 75t + 20b = 810 now, we know that the landscaping company policy is to plant 3 bushes for every tree so that makes the amount "b" 1 tree 1 bush + 1 bush + 1 bush 1 tree 1 bush + 1 bush + 1 bush 1 tree 1 bush + 1 bush + 1 bush so amount wise, if you say plant 3 trees, you'd plant 3(3) bushes that makes the bushes b = 3t so our equation will be \(\bf 75t + 20(3t) = 810\) then just solve for "t" and that'd give you how many trees, multiply that by 3, and that's how many bushes they planted